# Augmenting Subspace Optimization Methods with Linear Bandits

### v1

@@ 3.1 Understanding Theorem 1
Let the setting and assumptions of Theorem 2 hold. Suppose the iteration (3) is employed as the optimization algorithm in Framework 2 with a constant stepsize $\alpha_{k}=\frac{1}{L}$. Suppose that $U_{k}=\nabla f(x_{k})$ on every iteration of Algorithm 3. Then, with probability at least $1-\delta^{^{\prime}}$,
$D_{K}\leq\left(2M\displaystyle\sqrt{\frac{2dS_{\max}L(f(x_{0})-f(x_{*}))(M+1)}{q\lambda}}+\displaystyle\sqrt{\frac{16\lambda dS_{\max}L(f(x_{0})-f(x_{*}))}{qM}\log\left(1+\displaystyle\frac{M}{\lambda d}\right)}\right)\sqrt{K},$
@@
From Corollary 1 and Corollary 2, we see that, unsurprisingly, the regret bound is better when we have perfect knowledge of the magnitude $\|\nabla f(x_{k})\|$ on every iteration of Algorithm 3. This knowledge of $\|\nabla f(x_{k})\|$ obviously renders the algorithm analyzed in Corollary 2 impractical. Moreover, even the algorithm analyzed in Corollary 1 is impractical, although considerably less so; the algorithm considered in Corollary 1 requires knowledge of not only the gradient upper bound $G$ in 4 but also the global Lipschitz constant $L$ in 1. Regardless, Corollary 1 and Corollary 2 still suggest that, given a subspace optimization algorithm in Framework 1 that can yield upper bounds on $V_{K}$ similar to or betterthan the coarse bound from Theorem 2, the linear UCB mechanism in Algorithm 3 will exhibit sublinear dynamic regret, as we intended to demonstrate.
@@ 4 Practical Considerations
2(Inputs) Initial point $x_{1}$, horizon $K$, regularizer $\lambda>0$.
3(Initializations) Initial covariance matrix $C_{1}(\leftarrow\lambda I_{d})$, initial right-hand side $b_{1}(\leftarrow 0_{d})$, initial gradient estimate $g_{1}(\leftarrow 0_{d})$, memory parameter $M\geq 0$, average weighting parameter $\mu\in(0,1)$, backtracking parameter $\beta\in(0,1)$, sufficient decrease parameter $\sigma>0$, initial gradient upper bound estimate $U_{1}$.
@@
6      Determine/receive $S_{k}\in\mathbb{R}^{d\times p_{k}}$ and call oracle to compute $S_{k}^{\top}\nabla f(x_{k})$.
@@
9            $U_{1}\leftarrow\displaystyle\frac{d}{p_{1}}\|S_{1}^{\top}\nabla f(x_{1})\|$
@@
11            $U_{k}\leftarrow\mu U_{k-1}+(1-\mu)\displaystyle\frac{d}{p_{k}}\|S_{k}^{\top}\nabla f(x_{k})\|$
@@
14      Compute $s_{k}\leftarrow\displaystyle\arg\max_{s:\|s\|=1}g_{k}^{\top}s+\sqrt{\lambda}U_{k}\|s\|_{C_{k}^{-1}}$
@@
16      Call oracle to compute $r_{k}\leftarrow\langle\nabla f(x_{k}),s_{k}\rangle.$
@@
18      Compute $C_{k+1}$ and $b_{k+1}$ according to (12).
19      $g_{k+1}\leftarrow C_{k+1}^{-1}b_{k+1}$
20      (Generate trial step.)
21      $x_{+}\leftarrow x_{k}-\alpha_{k}P_{k}\nabla f(x_{k})$.
22      (Determine next incumbent and stepsize by checking sufficient decrease.)
23      if $f(x_{+})\leq f(x_{k})-\sigma\alpha_{k}\|P_{k}\nabla f(x_{k})\|^{2}$ then
24            $x_{k+1}\leftarrow x_{+}$
25            $\alpha_{k+1}=(1/\beta)\alpha_{k}$
26      else
27             $x_{k+1}\leftarrow x_{k}$
28            $\alpha_{k+1}\leftarrow\beta\alpha_{k}$
29
30
@@ 4.1 Incorporating arbitrary sketches
In particular, we suppose that given a sequence of sketch sizes $\{p_{k}\}_{k=1}^{K}$, we acquire a sequence of sketch matrices $\{S_{k}\in\mathbb{R}^{d\times p_{k}}\}$. After $s_{k}$ is computed in Equation 6, we augment $S_{k}$ by the column $s_{k}$ so that $S_{k}\in\mathbb{R}^{d\times(p_{k}+1)}$. We then add a second subscript to the variables $r$ and $s$ so that, for each iteration $k$, $\{r_{k,i}:i=1,\dots,p_{k},p_{k}+1\}$ corresponds to the entries of $S_{k}^{\top}\nabla f(x_{k})$ and and $\{s_{k_{i}}:i=1,\dots,p_{k},p_{k}+1\}$ corresponds to the columns of $S_{k}$. We then replace Algorithm 3 and Algorithm 3, respectively, with
$C_{k+1}\leftarrow\lambda I_{d}+\displaystyle\sum_{j=\max\{1,k-M\}}^{k}\displaystyle\sum_{i=1}^{p_{j}+1}s_{j,i}s_{j,i}^{\top}\quad\text{ and }\quad b_{k+1}\leftarrow\displaystyle\sum_{j=\max\{1,k-M\}}^{k}\displaystyle\sum_{i=1}^{p_{j}+1}r_{j,i}s_{j,i}.$ (12)
@@
$x_{k+1}\leftarrow x_{k}-\alpha_{k}P_{k}\nabla f(x_{k}),$
as in (3), noting that this incumbent update requires the additional (implicit) computation of $S_{k}(S_{k}^{\top}S_{k})^{-1}$.
4.2 Estimating $\|\nabla f(x_{k})\|$
We saw in comparing Corollary 2 to Corollary 1 that having knowledge of $\|\nabla f(x_{k})\|$ in each iteration $k$ for use in the subproblem (6) leads to better bounds on dynamic regret. While we generally cannot know $\|\nabla f(x_{k})\|$ in practice, we have found it is beneficial to approximate $\|\nabla f(x_{k})\|$ online. We accomplish this by a simple exponential moving average applied to the quantity $\|(S_{k}^{T}S_{k})^{-1}S_{k}^{\top}\nabla f(x_{k})\|$ computed in each iteration $k$; see Algorithm 4 of Algorithm 4.
4.3 Dynamic stepsizes
As in the framework proposed and analyzed in [9], our practical implementation does not fix a constant stepsize $\alpha_{k}$ but instead employs an adaptive stepsize that behaves similarly to a trust-region mechanism; see Algorithm 4 of Algorithm 4. This effectively overcomes the difficulty associated with needing to estimate $L$ in 1. We remark that the subspace optimization method being employed in Algorithm 4 amounts to exactly the quadratic regularization subspace optimization method of [9] with a model Hessian set to a matrix of all zeros.
@@ 4.4 Memory management
Throughout Algorithm 4 we manage memory by maintaining a dictionary of size $M$ of sketches $S_{k}$ and their corresponding linear measurements $r_{k}=S_{k}^{\top}\nabla f(x_{k})$. In the prototype implementation that we tested, the action of the inverse covariance matrix is performed by accessing the dictionary and iteratively applying the Sherman–Morrison–Woodbury formula.
@@ 4.5 A subspace variant of POUNDers
In addition to implementing Algorithm 4 as described, we applied the UCB mechanism discussed and analyzed in this manuscript to the DFO solver POUNDers [49] to yield a variant of POUNDers that fits within the subspace optimization Framework 1.
@@
As a model-based method, POUNDers maintains a history of previously evaluated points and their corresponding function evaluations, $\mathcal{Y}=\{((y^{1},f(y^{1})),\dots,(y^{|\mathcal{Y}|},f(y^{|\mathcal{Y}|}))\}$. POUNDers is, moreover, a model-based trust-region method. Common to all model-based trust-region DFO methods (see, e.g., [31][Section 2.2]), every iteration of POUNDers has both an incumbent $x_{k}$ and a trust-region radius $\Delta_{k}$. On every iteration, POUNDers applies Algorithm 5 to identify, given $x_{k}$, $\Delta_{k}$, and $\mathcal{Y}$, a set of mutually orthonormal columns $S$ and a perpendicular set of mutually orthonormal columns, $S^{\perp}$, so that a subset of points in $\mathcal{Y}$ both (1) are sufficiently close (measured as a factor of $\Delta)$ to $x_{k}$ and (2) exhibit sufficient volume in the affine subspace defined by $x_{k}$ and $S$ (as measured by projection onto the perpendicular subspace). Algorithm 5 is a greedy algorithm that grows the set of columns of $S$, one at a time, and is implemented by performing QR insertions.
@@
Having computed $S$ and $S^{\perp}$, POUNDers will then perform the function evaluations $\{f(x_{k}+\Delta_{k}p):p\text{ is a column of }S^{\perp}\}$, provided $S^{\perp}$ is nonempty; the points $\{x_{k}+\Delta_{k}p:p\text{ is a column of }S^{\perp}\}$ are referred to as geometry points. With these geometry points now included in $\mathcal{Y}$, one can prove that there exists a subset of $\mathcal{Y}$ so that a quadratic model interpolating those points locally (as measured by $\Delta_{k}$) exhibits errors of the same order as a first-order Taylor model; see, for example, [48][Theorem 4.1].
@@
In this aspect of POUNDers we make a large modification to yield “subspace POUNDers,” or SS-POUNDers for short. Rather than evaluating a stencil in $S^{\perp}$ to guarantee local accuracy of a model of $f$ in the full space $\mathbb{R}^{d}$ , SS-POUNDers employs a confidence ellipse $C_{k}$ and mean estimate of a gradient $g_{k}$, as in Algorithm 3, and it selects a single column $s_{k}$ of $S^{\perp}$ by solving
$s_{k}\leftarrow\arg\max_{s\in S^{\perp}}g_{k}^{\top}s+\sqrt{\lambda}U_{k}\|s\|_{C_{k}^{-1}}.$ (13)
@@
In SS-POUNDers, we then augment $S$ from Algorithm 5 by the column $s_{k}$ to yield $S_{k}$. We can then augment $S_{k}$ by additional randomized sketches in the span of $S^{\perp}$; this additional randomization is performed identically to [12][Algorithm 5], and we provide pseudocode in Algorithm 7. We then construct a quadratic interpolation model $m_{k}:\mathbb{R}^{dim(span(S_{k}))}\to\mathbb{R}$ using only points in the affine subspace defined by $x_{k}$ and $S_{k}$; this is accomplished by using the same model-building techniques as in POUNDers but by fitting the quadratic model to projected points $S_{k}^{\top}(y-x_{k})$ for the selected subset of points $y\in\mathcal{Y}$. We then minimize this lower-dimensional model over a trust-region to obtain an (approximate) minimizer $z_{k}\in\mathbb{R}^{dim(span(S_{k}))}$, which is embedded back into the full space via the transformation $x_{k}+S_{k}z_{k}$. This idea of minimizing a derivative-free model only over a selected affine subspace and then re-embedding the solution in the full space is the same idea employed in, for instance, [12]. SS-POUNDers updates the estimators $C_{k}$ and $b_{k}$ using both $S_{k}$ and the subspace derivative-free model gradient. The rest of SS-POUNDers appears essentially the same as POUNDers, and pseudocode is provided in Algorithm 6.
@@
1 Input: Center point $x_{k}\in\mathbb{R}^{d}$, bank of evaluated points $\mathcal{Y}=\{(y^{1},f(y^{1}),\dots,(y^{|\mathcal{Y}|},f(y^{|\mathcal{Y}|}))\}$ satisfying $(x,f(x))\in\mathcal{Y}$, trust-region radius $\Delta_{k}$.
@@
3 Set $S=\{s^{1}\}=\{0_{d}\}$.
4 Set $S^{\perp}=I_{d}$.
@@
6       if $\|y^{i}-x\|\leq c\Delta_{k}$ and $\left|\text{proj}_{S^{\perp}}\left(\frac{1}{c\Delta_{k}}(y^{i}-x)\right)\right|\geq\theta_{1}$ then
7             $S=S\cup\{y^{i}-x\}$
8             Update $S^{\perp}$ to be an orthonormal basis for $\mathcal{N}([s^{2}\cdots s^{|S|}])$
@@ 7             $S=S\cup\{y^{i}-x\}$
10      if $|S|=n+1$ then
11             break (the for loop)
@@
14 $Q=[S\hskip 6.0ptS^{\perp}]$
15 Return: $S,S^{\perp},Q$

### v4

@@ 3.1 Understanding Theorem 1
Let the setting and assumptions of Theorem 2 hold. Suppose Algorithm 2 is employed as the optimization algorithm in Framework 2. Suppose that $U_{k}=\|\nabla f(\mathbf{x}_{k})\|$ on every iteration of Algorithm 4. Then, with probability at least $1-\delta^{^{\prime}}$, $D_{K}\leq C\sqrt{K}$, where
$C:=2M\displaystyle\sqrt{\frac{2dS_{\max}L(f(\mathbf{x}_{0})-f(\mathbf{x}_{*}))(M+1)}{q(2\beta-1)\lambda}}+\displaystyle\sqrt{\frac{16\lambda dS_{\max}L(f(\mathbf{x}_{0})-f(\mathbf{x}_{*}))}{q(2\beta-1)M}\log\left(1+\displaystyle\frac{M}{\lambda d}\right)}$
@@
From Corollary 1 and Corollary 2, we see that, unsurprisingly, the regret bound is better when we have perfect knowledge of the magnitude $\|\nabla f(\mathbf{x}_{k})\|$ on every iteration of Algorithm 4. This knowledge of $\|\nabla f(\mathbf{x}_{k})\|$ obviously renders the algorithm analyzed in Corollary 2 impractical. Moreover, even the algorithm analyzed in Corollary 1 is impractical, although considerably less so; the algorithm considered in Corollary 1 requires knowledge of the gradient upper bound $G$ in 4. Regardless, Corollary 1 and Corollary 2 still suggest that, given a subspace optimization algorithm in Framework 1 that can yield upper bounds on $V_{K}$ similar to or better than the coarse bound from Theorem 2, the linear UCB mechanism in Algorithm 4 will exhibit sublinear dynamic regret, as we intended to demonstrate.
@@ 4 Practical Considerations
2(Inputs) Initial point $\mathbf{x}_{1}$, horizon $K$, regularizer $\lambda>0$.
3(Initializations) Initial covariance matrix $\mathbf{C}_{1}(\leftarrow\lambda\mathbf{I}_{d})$, initial right-hand side $\mathbf{b}_{1}(\leftarrow\mathbf{0}_{d})$, initial gradient estimate $\mathbf{g}_{1}(\leftarrow\mathbf{0}_{d})$, memory parameter $M\geq 0$, average weighting parameter $\mu\in(0,1)$, backtracking parameter $\beta\in(0,1)$, sufficient decrease parameter $\sigma>0$, initial gradient upper bound estimate $U_{1}$.
@@
6  Receive $\mathbf{S}_{k}\in\mathbb{R}^{d\times p_{k}}$ and call oracle to compute $\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})$.
@@
9    $U_{1}\leftarrow\displaystyle\frac{d}{p_{1}}\|\mathbf{S}_{1}^{\top}\nabla f(\mathbf{x}_{1})\|$.
@@
11    $U_{k}\leftarrow\mu U_{k-1}+(1-\mu)\displaystyle\frac{d}{p_{k}}\|\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|$.
@@
14  Compute $\mathbf{s}_{k}\leftarrow\displaystyle\arg\max_{\mathbf{s}:\|\mathbf{s}\|=1}\mathbf{g}_{k}^{\top}\mathbf{s}+\sqrt{\lambda}U_{k}\|\mathbf{s}\|_{\mathbf{C}_{k}^{-1}}$.
@@
16  Call oracle to compute $r_{k}\leftarrow\langle\nabla f(\mathbf{x}_{k}),\mathbf{s}_{k}\rangle.$
@@
18  Compute $\mathbf{C}_{k+1}$ and $\mathbf{b}_{k+1}$ according to (11).
19 $\mathbf{g}_{k+1}\leftarrow\mathbf{C}_{k+1}^{-1}\mathbf{b}_{k+1}$.
20 (Determine next incumbent by backtracking line search.)
21 $\alpha\leftarrow 1$.
22 $\mathbf{x}_{+}\leftarrow\mathbf{x}_{k}-\alpha\mathbf{P}_{k}\nabla f(\mathbf{x}_{k})$.
23 while $f(\mathbf{x}_{+})\geq f(\mathbf{x}_{k})-\sigma\alpha_{k}\|\mathbf{P}_{k}\nabla f(\mathbf{x}_{k})\|^{2}$ do
24    $\alpha\leftarrow\beta\alpha.$
25    $\mathbf{x}_{+}\leftarrow\mathbf{x}_{k}-\alpha\mathbf{P}_{k}\nabla f(\mathbf{x}_{k})$.
26
27 $\mathbf{x}_{k+1}\leftarrow\mathbf{x}_{+}$.
28
@@ 4.1 Incorporating arbitrary sketches
In particular, we suppose that given a sequence of sketch sizes $\{p_{k}\}_{k=1}^{K}$, we acquire a sequence of sketch matrices $\{\mathbf{S}_{k}\in\mathbb{R}^{d\times p_{k}}\}$. After $\mathbf{s}_{k}$ is computed in Algorithm 4, we augment $\mathbf{S}_{k}$ by the column $\mathbf{s}_{k}$ so that $\mathbf{S}_{k}\in\mathbb{R}^{d\times(p_{k}+1)}$. We then add a second subscript to the variables $r$ and $\mathbf{s}$ so that, for each iteration $k$, $\{r_{k,i}:i=1,\dots,p_{k},p_{k}+1\}$ corresponds to the entries of $\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})$ and and $\{\mathbf{s}_{k_{i}}:i=1,\dots,p_{k},p_{k}+1\}$ corresponds to the columns of $\mathbf{S}_{k}$. We then replace Algorithm 4 and Algorithm 4, respectively, with
$\mathbf{C}_{k+1}\leftarrow\lambda\mathbf{I}_{d}+\displaystyle\sum_{j=\max\{1,k-M\}}^{k}\displaystyle\sum_{i=1}^{p_{j}+1}\mathbf{s}_{j,i}\mathbf{s}_{j,i}^{\top}\quad\text{ and }\quad\mathbf{b}_{k+1}\leftarrow\displaystyle\sum_{j=\max\{1,k-M\}}^{k}\displaystyle\sum_{i=1}^{p_{j}+1}r_{j,i}\mathbf{s}_{j,i}.$ (11)
@@
$\mathbf{x}_{k+1}\leftarrow\mathbf{x}_{k}-\alpha_{k}\mathbf{P}_{k}\nabla f(\mathbf{x}_{k}),$
as in Algorithm 2, noting that this incumbent update requires the additional (implicit) computation of $\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}$.
4.2 Estimating $\|\nabla f(\mathbf{x}_{k})\|$
We saw in comparing Corollary 2 to Corollary 1 that having knowledge of $\|\nabla f(\mathbf{x}_{k})\|$ in each iteration $k$ for use in the subproblem (5) leads to better bounds on dynamic regret. While we generally cannot know $\|\nabla f(\mathbf{x}_{k})\|$ in practice, we have found it is beneficial to approximate $\|\nabla f(\mathbf{x}_{k})\|$ dynamically. We accomplish this by a simple exponential moving average applied to the quantity $\|(\mathbf{S}_{k}^{T}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|$ computed in each iteration $k$; see Algorithm 5 of Algorithm 5.
@@ 4.3 Memory management
Throughout Algorithm 5 we manage memory by maintaining a dictionary of size $M$ of sketches $\mathbf{S}_{k}$ and their corresponding linear measurements $r_{k}=\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})$. In the prototype implementation that we tested, the action of the inverse covariance matrix is performed by accessing the dictionary and iteratively applying the Sherman–Morrison–Woodbury formula; see, e.g., [27, Equation 2].
@@ 4.4 A subspace variant of POUNDers
In addition to implementing Algorithm 5 as described, we applied the UCB mechanism discussed and analyzed in this manuscript to the DFO solver POUNDers [56] to yield a variant of POUNDers that fits within the subspace optimization Framework 1. Our motivation for introducing this variant is to demonstrate that the application of the linear bandit method of Algorithm 4 to any method in Framework 1 is relatively straight-forward, and need not only apply to the fairly simple Algorithm 2.
@@
As a model-based method, POUNDers maintains a history of previously evaluated points and their corresponding function evaluations, $\mathcal{Y}=~\{((\mathbf{y}^{1},f(\mathbf{y}^{1})),\dots,(\mathbf{y}^{|\mathcal{Y}|},f(\mathbf{y}^{|\mathcal{Y}|}))\}$. POUNDers is, moreover, a model-based trust-region method. Common to all model-based trust-region DFO methods (see, e.g., [36, Section 2.2]), every iteration of POUNDers has both an incumbent $\mathbf{x}_{k}$ and a trust-region radius $\Delta_{k}$. On every iteration, POUNDers applies Algorithm 6 to identify, given $\mathbf{x}_{k}$, $\Delta_{k}$, and $\mathcal{Y}$, a set of mutually orthonormal columns $\mathbf{S}$ and a perpendicular set of mutually orthonormal columns, $\mathbf{S}^{\perp}$, so that a subset of points in $\mathcal{Y}$ both (1) are sufficiently close (measured as a factor of $\Delta)$ to $\mathbf{x}_{k}$ and (2) exhibit sufficient volume in the affine subspace defined by $\mathbf{x}_{k}$ and $\mathbf{S}$ (as measured by projection onto the perpendicular subspace). Algorithm 6 is a greedy algorithm that grows the set of columns of $\mathbf{S}$, one at a time, and is implemented by performing QR insertions.
@@
Having computed $\mathbf{S}$ and $\mathbf{S}^{\perp}$, POUNDers will then perform the function evaluations $\{f(\mathbf{x}_{k}+\Delta_{k}\mathbf{p}):\mathbf{p}\text{ is a column of }\mathbf{S}^{\perp}\}$, provided $\mathbf{S}^{\perp}$ is nonempty; the points $\{\mathbf{x}_{k}+~\Delta_{k}\mathbf{p}:\mathbf{p}\text{ is a column of }\mathbf{S}^{\perp}\}$ are referred to as geometry points. With these geometry points now included in $\mathcal{Y}$, one can prove that there exists a subset of $\mathcal{Y}$ so that a quadratic model interpolating those points locally (as measured by $\Delta_{k}$) exhibits errors of the same order as a first-order Taylor model; see, for example, [55, Theorem 4.1].
@@
In this aspect of POUNDers we make a large modification to yield “subspace POUNDers,” or SS-POUNDers for short. Rather than evaluating a stencil in $\mathbf{S}^{\perp}$ to guarantee local accuracy of a model of $f$ in the full space $\mathbb{R}^{d}$ , SS-POUNDers employs a confidence ellipse $\mathbf{C}_{k}$ and mean estimate of a gradient $\mathbf{g}_{k}$, as in Algorithm 4, and it selects a single column $\mathbf{s}_{k}$ of $\mathbf{S}^{\perp}$ by solving
$\mathbf{s}_{k}\leftarrow\arg\max_{\mathbf{s}\in\mathbf{S}^{\perp}}\mathbf{g}_{k}^{\top}\mathbf{s}+\sqrt{\lambda}U_{k}\|\mathbf{s}\|_{\mathbf{C}_{k}^{-1}}.$ (12)
@@
In SS-POUNDers, we then augment $\mathbf{S}$ from Algorithm 6 by the column $\mathbf{s}_{k}$ to yield $\mathbf{S}_{k}$. We can then augment $\mathbf{S}_{k}$ by additional randomized sketches in the span of $\mathbf{S}^{\perp}$; this additional randomization is performed identically to [12, Algorithm 5], and we provide pseudocode in Algorithm 8. We then construct a quadratic interpolation model $m_{k}:\mathbb{R}^{\mathrm{dim}(\mathrm{Span}(\mathbf{S}_{k}))}\to\mathbb{R}$ using only points in the affine subspace defined by $\mathbf{x}_{k}$ and $\mathbf{S}_{k}$; this is accomplished by using the same model-building techniques as in POUNDers but by fitting the quadratic model to projected points $\mathbf{S}_{k}^{\top}(\mathbf{y}-\mathbf{x}_{k})$ for the selected subset of points $\mathbf{y}\in\mathcal{Y}$. We then minimize this lower-dimensional model over a trust-region to obtain an (approximate) minimizer $\mathbf{z}_{k}\in\mathbb{R}^{\mathrm{dim}(\mathrm{Span}(\mathbf{S}_{k}))}$, which is embedded back into the full space via the transformation $\mathbf{x}_{k}+\mathbf{S}_{k}\mathbf{z}_{k}$. This idea of minimizing a derivative-free model only over a selected affine subspace and then re-embedding the solution in the full space is the same idea employed in, for instance, [12]. SS-POUNDers updates the estimators $\mathbf{C}_{k}$ and $\mathbf{b}_{k}$ using both $\mathbf{S}_{k}$ and the subspace derivative-free model gradient. The rest of SS-POUNDers appears essentially the same as POUNDers, and pseudocode is provided in Algorithm 7.
@@
1Input: Center point $\mathbf{x}_{k}\in\mathbb{R}^{d}$, bank of evaluated points $\mathcal{Y}=\{(\mathbf{y}^{1},f(\mathbf{y}^{1}),\dots,(\mathbf{y}^{|\mathcal{Y}|},f(\mathbf{y}^{|\mathcal{Y}|}))\}$ satisfying $(\mathbf{x}_{k},f(\mathbf{x}_{k}))\in\mathcal{Y}$, trust-region radius $\Delta_{k}$.
@@
3 Set $\mathbf{S}=\{\mathbf{s}^{1}\}=\{\mathbf{0}_{d}\}$.
4 Set $\mathbf{S}^{\perp}=\mathbf{I}_{d}$.
@@
6 if $\|\mathbf{y}^{i}-\mathbf{x}\|\leq c\Delta_{k}$ and $\left|\text{proj}_{\mathbf{S}^{\perp}}\left(\frac{1}{c\Delta_{k}}(\mathbf{y}^{i}-\mathbf{x})\right)\right|\geq\theta_{1}$ then
7    $\mathbf{S}=\mathbf{S}\cup\{\mathbf{y}^{i}-\mathbf{x}\}$.
8      Update $\mathbf{S}^{\perp}$ to be an orthonormal basis for $\mathrm{Null}([\mathbf{s}^{2}\cdots\mathbf{s}^{|S|}])$.
@@
10 if $|\mathbf{S}|=n+1$ then
11    break (the for loop).
@@
14 $\mathbf{Q}=[\mathbf{S}\hskip 6.0pt\mathbf{S}^{\perp}]$.
15 Return: $\mathbf{S},\mathbf{S}^{\perp},\mathbf{Q}$